In a closed economy with no govt Consumption function is C=100+0.8Y where Y is income
What happens to income Y if investors become more optimistic snd increase investment by 20?
Looking for an explanation with the answer as to Y moves in the direction it does.
1
Expert's answer
2019-02-21T10:50:08-0500

When the economy is in equilibrium

GDP = C + I + G + NX,

where

C- consumption,

I – investment,

G – government spending,

NX – net exports.

We know that C=100+0.8Y

ΔI = 20

G=0

NX=0, because the economy is closed.

Y=100+0.8Y+20+0+0

0.2Y=120

Y=600

The income will increase by 600.